OpenStudy (anonymous):

How do we convert this function into a logarithmic one? P(x) = 100(1 + 0.1)^x

OpenStudy (aonz):

Well we could use magic but lets not :P Anyways! rewrite to be p(x) / 100 = (1+0.1)^x Remember that If y = 2^x x = log(2) y

OpenStudy (anonymous):

I can isolate to P(x) /100 = 100*(1.1)^x / 100 P(x) /100 = (1.1)^x but then I dont understand how a function can be P(x)/100 If I just change it to y I get y /100 = (1.1)^x and then I use the log conversion y = b^x ↔ x = log_b y and get x = log_1.1 (y/100) but now how do I get back to a function ? like how do I isolate y again? I think I need something like... P(x) = ??? If I try to use something like the quotient property x = log_1.1 (y) - log_1.1 (100) then maybe x +log_1.1 (100) = log_1.1 (y) - log_1.1 (100) + log_1.1 (100) x +log_1.1 (100) = log_1.1 (y) umm then somehow try to convert log_1.1 (y) into something that lets me isolate y if y = b^x ↔ x = log_b y then I guess y =1.1^(x +log_1.1 (100) ) And so P(x) = 1.1^(x +log_1.1 (100) ) That all looks very wrong.. or is that actually right in some way? am I confused enough? What am I missing here.

OpenStudy (aonz):

Okayy. I need some time to read this :P The question is simply rewriting to log function right?

OpenStudy (anonymous):

Actually,maybe it is just that I don't understand what they mean by "convert your exponential function P(x) into a logarithmic one" Does this mean that while P(x) originally would return the product after applying an exponent (x), that now they want to return the exponent after giving it a product so x=a target balance for example and it returns a number of years? Is that what they mean by a logarithmic function?

OpenStudy (aonz):

Looking at the question, this should be fine.